Problem 10.30 A 21.5 mL sample of a KOH solution o
Answer #1

The reaction that takes place is:
H2SO4 + 2 KOH ----> K2SO4 + 2H2O

from given reaction for neutralisation,
2*number of moles of H2SO4 = number of moles of KOH
2*M(H2SO4)*V(H2SO4) = M(KOH) * V(KOH)
2*0.250*11.1 = M*21.5
M = 0.258 M

Answer: 0.258 M